OpenStudy (anonymous):

State whether the given measurements determine zero, one, or two triangles. A = 61°, a = 23, b = 24

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

@bahrom7893 ?

OpenStudy (anonymous):

@RadEn

OpenStudy (bahrom7893):

i can't think of those triangle identities.. satellite help us out, it's been ages since i've done geometry lol

OpenStudy (bahrom7893):

I know there's a theorem that involves an angle and two sides

OpenStudy (anonymous):

@satellite73

OpenStudy (bahrom7893):

ok let's research this a bit and try this

OpenStudy (anonymous):

@terenzreignz

OpenStudy (anonymous):

@radar

OpenStudy (bahrom7893):

Let's use the theory of sins

OpenStudy (bahrom7893):

SinA/a = SinB/b = SinC/c So in your case: Sin61/23 = SinB/24

OpenStudy (bahrom7893):

SinB = (24*Sin61)/23 = approximately 0.912

OpenStudy (anonymous):

So there is only one?

OpenStudy (bahrom7893):

Sorry was trying to find my calculator.. anyway, now if SinB = 0.912 B = ArcSin(0.912) which is about 65.78 degrees

OpenStudy (bahrom7893):

This means that we have the following angles: A = 61 B = 66 C = 180 - 61 - 66 = 53

OpenStudy (bahrom7893):

So i guess only 1